A 415V to 11 kV transformer has a rating of 200 kVA. The winding
resistance and leakage reactance when referred to the primary are 0.014 ?
and 0.057 ohms respectively.

(a) Determine the % regulation of the transformer at 0.8 power factor lagging.

(b) In designing a particular 415V to 11 kV, 200 kVA transformer, the
primary winding resistance is to be 10 m?. Find the maximum winding
resistance of the secondary winding if the transformer is to have 2%
regulation at unity power factor
1
Expert's answer
2011-03-17T07:03:22-0400
v(t) = int(a(t), dt) + v(0) = sin(t) - cos(t) +v(0) = sin(t) - cos(t) + 3
s(t) = int(v(t), dt) + s(0) = -cos(t) - sin(t) + 3t + 3